OpenStudy (anonymous):

Find the value of the permutation P(8,3)

OpenStudy (zarkon):

336

OpenStudy (anonymous):

Lol, Zarkon really? :D

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

The answer is right, but this is just too easy for Zarkon :)

OpenStudy (zarkon):

8*7*6...feeling lazy...just looking at easy problems

OpenStudy (anonymous):

can you tell me what the formula for permutation is?

OpenStudy (anonymous):

\[P(n,r)= n \times (n-1) \times (n-2) \times \cdots \times (n-r+1)\]

OpenStudy (anonymous):

thank you!